# TikTokActions: A TikTok-Derived Video Dataset for Human Action Recognition

### v2

@@ Acknowledgements
This work was made possible by a multitude of supporters and funding sources. We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No. 2118222) under the HI-DSI project. The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services – Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable. The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y. Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J. Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: https://doi.org/10.1145/3437359.3465565). The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work. We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858).
@@ Appendix C Dataset Composition
Figure 5: Pie chart categorization of video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.

### v3

@@ Acknowledgements
This work was made possible by a multitude of supporters and funding sources. We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No. 2118222) under the HI-DSI project. The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services–Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable. The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y. Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J. Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: 10.1145/3437359.3465565). The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work. We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858). We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly. However, all contents in this manuscript are original ideas and analyses conducted by the authors.
@@ Appendix C Dataset Composition
Figure 5: Video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.